OpenStudy (anonymous):

Please help with this easy question! I need to find x and y from these 2 equations: e^-y * 2x = 0 2ye^-y - e^-y(y^2) = 0 I already found x=0 by solving for x in the first equation. and then i plugged in 0 for x into the 2nd equation and I simplified the equation to ye^-y(2-ye^-y)=0 but im stuck here.. @amistre64

OpenStudy (anonymous):

oh it has to be both?

OpenStudy (anonymous):

there is no \(x\) in the second equation

OpenStudy (anonymous):

@satellite73 im sorry, 2nd equation: 2ye^-y - e^-y(x^2+y^2)=0. I plugged in 0 for x and I simplified it to ye^-y(2-ye^-y)=0

OpenStudy (anonymous):

Im trying to get the critical points right now and I need to find x and y from these 2 equations that I obtained by taking the partial derivatives

OpenStudy (anonymous):

so you end up with \[2ye^{-y}-y^2e^{-y}=0\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then i factored out ye^-y

OpenStudy (anonymous):

ok then factor out the \(e^{-y}\) or just ignore it because it is never zero, and solve \[y-y^2=0\]

OpenStudy (anonymous):

oops rather \[2y-y^2=0\]

OpenStudy (anonymous):

should be real easy now right?

OpenStudy (anonymous):

Yes thank you!

OpenStudy (anonymous):

yw